A student went to test the motorbike one afternoon after class on the Parys Road. His velocity between two hills is listed in the table. The time is in seconds and the speed readings have been converted to m/s to ease calculations.

t(s) | 1 | 5 | 7 |

v(m/s)| 24 | 40 | 36 |


a) Determine the Lagrange functions L0, L1 and L2 and the second degree Lagrange interpolation polynomial P2(t) for the given data points. Simplify the polynomial.


b) formulate a theorem on the existence of a unique interpolation polynomial.




1
Expert's answer
2021-11-09T01:09:06-0500

The given values are x0 = 0, y0 = -12 x1 = 1, y1 = 0 x2 = 3, y2 = 6 x3 = 4, y3 = 12 By Lagrange’s interpolaiton formula, ⇒ y = (x – 1) (x – 3) (x – 4) – x (x – 1) (x – 4) + x (x – 1) (x – 3) ⇒ y = (x – 1) (x – 4)[(x – 3) – x] + x (x – 1) (x – 3) ⇒ y = (x – 1) (x – 4) (-3) + x (x – 1) (x – 3) ⇒ y = (x – 1) [-3x + 12 + x2 – 3x] ⇒ y = (x – 1) (x2 – 6x + 12) ⇒ y = x3 – 6x2 + 12x – x2 + 6x – 12 y = x3 – 7x + 18x – 12 is the required polynomial formulae